OpenStudy (anonymous):

Tomas wrote the equation y = 3x +. When Sandra wrote her equation, they discovered that her equation had all the same solutions as Tomas’s equation. Which equation could be Sandra’s?

OpenStudy (anonymous):

–6x + y = 6x + y = –6x + 2y = 6x + 2y =